# Surveying the landscape of optically addressable spin qubits for quantum information and sensing technology

### v1

@@ Abstract
Quantum technologies offer ways to solve certain tasks more quickly, efficiently, and with greater sensitivity than their classical counterparts. Yet substantial challenges remain in the construction of sufficiently error-free and scaleable quantum platforms that are needed to unlock any real benefits to society. Acknowledging that this hardware can take vastly different forms, our review here focuses on so-called spintronic (i.e. spin-electronic) materials that use electronic or nuclear spins to embody qubits. Towards helping the reader to spot trends and pick winners, we have surveyed the various families of optically addressable spin qubits and attempted to benchmark and identify the most promising ones in each group. We reveal further trends that demonstrate how qubit lifetimes depend on the material’s synthesis, the concentration/distribution of its embedded qubits, and the experimental conditions.
Keywords electron spin qubits quantum sensing spintronics materials engineering
@@ 1 Introduction
Over the last 40 years, quantum computation has progressed from concept to hardware. In 2023, IBM reported the commissioning of a 1121-qubit superconducting quantum processor[1], while Honeywell demonstrated a quantum charge-coupled device architecture based on trapped-ions[2]. Despite these achievements, the scalability and thus ultimate utility of each species of hardware remains highly contested.
@@
Recently, spintronic (spin electronic) materials have garnered attention as contenders for qubit media. The best of these materials offer coherence times exceeding milliseconds, albeit at cryogenic temperatures[3]. Progress in the field has made the dream of usefully storing quantum information and/or implementing quantum processing at scale with appropriately interacting spins more tangible. Start-up company Quantum Brilliance[4], recently demonstrated a room-temperature (RT) qubit system based on nitrogen-vacancy (NV) centres in diamond, for example. Beyond computation, these same spintronic materials are already providing advantageous forms of sensing, especially under ambient conditions[5]. This literature review focuses, from a materials science perspective, on optimising DiVincenzo’s criterion, namely achieving long effective qubit decoherence in solid-state spin-based systems.
@@
Acknowledging that multiple levels (qudits) can be used advantageously for error correction, quantum information is most simply encoded as the state of a two-level system. When necessary, a magnetic or electric field is applied to split otherwise degenerate sublevels. The higher and lower energy sublevels, and , form two orthogonal basis states. A pure state is formed as a coherent superposition of the two (), and the loss of this coherence results in information loss (and thus errors). For convenience, we visualise pure states as points on the surface of a Bloch sphere (Figure 2), each located by a polar angle and an azimuthal angle [6]. The value of determines the probability amplitude of finding the qubit in either the or state, while determines the relative phase of the qubit between the and states. Our ability to manipulate and exploit the quantum state of a material is ultimately limited by the lifetime of any given state, which measured by, depending on the precise application, three different time parameters, namely, the spin-lattice relaxation time, , the spin coherence time, , (also known as phase memory time, ), and the spin-dephasing time (or free induction decay time), .
1.1 Important Quantum Spin Parameters
represents the time it takes for the "longitudinal magnitude" of the state vector to decay by a factor of e. is the time required for the transverse component of the system’s spin state vector to decay by a factor of e. Put simply by DiVincenzo[7], decoherence characterises the interactions of a qubit with its environment, causing a qubit in a generic (pure) state to relax to a mixed state. This causes the loss of information in a quantum system via the loss of phase coherence. is related to the homogeneous emission linewidth, f, by 1/. Finally, is a measure of the spin coherence time that includes the effects of inhomogeneous broadening due to fluctuations in local magnetic fields, B, within spin material. According to Chavhan et al., the relationship between and can be represented as 1/ = 1/ + B, where is the gyromagnetic ratio[8].
The methods used to determine , , and encompass CW and pulsed EPR[9], optically-detected magnetic resonance (ODMR)[10], and electrically detected magnetic resonance (EDMR)[11, 12]. A pulse sequence can be considered a filter that partially eliminates noise from the surroundings of a qubit system (Figure 2). Inversion and saturation recovery are common pulse sequences for measurements. Saturation recovery involves applying a strong and long pulse or a series of short pulses to achieve the equipartition of excited and ground states. Because this pulse sequence is long, the effects of spectral diffusion average out, hence the measurement closer tot the "intrinsic" can be achieved. On the other hand, inversion recovery only uses a short pulse to "flip" spins. This results in a measurement affected by spectral diffusion, hence a shorter [13]. Referring to Figure 2, the measurement sequence of only applies pulses to transform the state onto the transverse plane (-plane) and back to the longitudinal axis (z-axis). This allows the dephasing of a qubit to evolve without correction.
In general, sets an upper limit on and, since energy relaxation proceeds more slowly than decoherence, is (often substantially) longer than [14]. Likewise, sets an upper limit on (the pulse sequence used to measure deliberately attempts to reduce the effects of inhomogeneity). Thus, in general, .
@@ 2 Overview of Candidate Spin Qubit Materials
Quantum systems capable of operating at room temperature open up many additional applications that the overhead of cryogenic operation precludes. But, achieving high fidelity (in the initialisation, gate operations and read-out) of room-temperature qubits remains extremely challenging. In this review, we have attempted to survey the available "fully optically addressable materials" (FOAMs) that are most directly relevant to quantum information and sensing technology. We do not provide an exhaustive account of all reported spin systems. Rather, we present the different approaches, materials wise, that have shown promise and report the properties of the best-performing representatives of each approach.
FOAMs are an attractive option for quantum applications since they typically involve spin-levels that are energetically separated by far more than kBT and can be initialised using visible/near-infrared light to generate quasi-"pure" quantum states. Furthermore, optical signals, being themselves far larger than kBT, are less affected by thermal noise and benefit from excellent single-photon detectors that are available even at RT. Since the energies of the photons involved in both processes greatly exceed kBT, room-temperature operation at high levels of fidelity is possible. Such materials are usually probed using ODMR spectroscopy, where the quantum sensing sensitivity for a.c. magnetic fields, , is proportional to , where is duration of the readout process, is the measurement contrast, is the number of spins and is the number of scan averages. Relatively few FOAMs have been demonstrated so far and are often limited by their photoluminescence yields or relatively short quantum spin properties. For qubit media, Weber2010 et al., developed criterion for optically addressable solid-state spin defects[15], denoting that:
@@
A well-established example of a FOAM which satisfies these criteria is the negatively charged nitrogen-vacancy centres in diamond (herein simply, NV-diamond), which utilise the triplet ground state and hyperfine splitting arising from the NV-centre. Initialisation, (i.e., electron spin polarisation) is achieved by optical excitation with green light. The newly generated excited states either relax by emission of 637 nm light or undergo intersystem crossing (ISC) into a metastable singlet state (Figure 3a). Repopulation of the ground state then follows an intermediate relaxation between two singlet states (with emission at 1042 nm) and finally spin-selective ISC into the T0 sublevel, resulting in a strong spin polarisation. At zero-applied magnetic field (ZF) the T±1 states are degenerate due to the defects C3V symmetry. The application of a magnetic field lifts this degeneracy which enables the manipulation of spin density by microwave pulses thereby modify the fluorescence at characteristic Zeeman splitting frequencies (), which can be detected following futher optical excitations.
3D spin-defect FOAMs
@@
NV-diamond has been a cornerstone of ODMR-based quantum sensing due to its robust spin properties even at RT. Due to the relatively low spin densities (resulting in small dipolar coupling between spins) and the mismatch between diamond lattice vibrations (phonons) and the Larmor frequency of electrons spins in NV-’s S=1 ground state, NV-diamond can exhibit s of several milliseconds at RT[16, 17]. The nature and mechanism of its spin-lattice relaxation as a function of temperature, which includes an ‘Orbach-like" process (dependent on the phonon density at the spin transition frequency) and spin-phonon Raman scattering (of either first [18] or else second [17] order) scaling like , have been repeatedly investigated. The material’s quantum spin properties are highly dependent on the NV-centre depth[19], concentration[20], crystal strain[21], and the presence of impurities such as , N-centres or EPR-inactive neutral or positively charged NV-centres[22, 23]. Its popularity has seen a plethora of investigations to understand and modulate its spin properties. For example, high-field EPR spectroscopy and temperature-dependent measurements have demonstrated that decoherence from the N and 13C flip-flop fluctuations can be almost eliminated at low temperatures where the spin-bath is polarised[24]. Here, reached 250 s at 2 K following a sharp increase below 12 K in HTHP diamond samples. Achieving similar properties under low-field and higher temperature conditions with thermally polarised nuclei is challenging and requires meticulous materials preparation and/or the use of dynamical decoupling methods (vide infra). For example, the impact of parasitic nuclear spins and impurities was shown most remarkably by Balasuramanian et al.,[25]. Careful growth by CVD on a diamond substrate using isotopically enriched feedstock led to just 0.3 13C abundance and low levels of other paramagnetic impurity resulting in 1.8 ms at RT.
@@
More recently, increasing attention has been paid to using diamonds as a host of other spin defects owing to its wide band gap, efficient thermal dissipation, and physical and chemical stability. As such, various magnetically-active colour centres have been investigated for quantum applications[26]. These include HV[27], BV[28], OV[29], so-called group-IV vacancy defects[30] such as SiV[31, 32], SnV[33, 34], GeV[35], and PbV[36, 37], and even transition metal defects originating as impurities like NiV[38]. To our knowledge, spin polarisation has not been observed in HV or OV centres, whilst ODMR experiments on BV and PbV have not been reported.
These metal spin-centres typically exhibit strong spin-orbit coupling (several hundred GHz) that gives rise to zero-field splitting (ZFS) even for S=1/2 species, as well as electron-nuclear spin coupling making them qudit candidates. Moreover, group-IV defects otherwise exhibit useful photonic properties that make them attractive for quantum applications such as Fourier-limited ZPL linewidths, high spectral stability, coherent photon emission, and strain-responsive band gap engineering[39]. For example, negatively charged SiV (S=1/2 system) has enabled direct observation of photon interference[40]. Outside of a dilution refrigerator, the spin coherence lifetimes are severely impacted by thermal acoustic phonon coupling. At 100 mK, reaches 1 second while the longest measured was 1.5 s. Spin coherence could be maintained by dynamical decoupling up to 600 mK where measured 60 s. can also be improved through strain engineering which modifies the spin-orbit coupling ground state splitting and subsequently the spin-phonon coupling[41]. By comparison, the neutral SiV (S=1 ground state) demonstrates an impressive 25 s, 0.1 ms even at 15 K, decreasing to 7.8 and 2 s at RT [42]. However, a route to reliably synthesising SiVs in diamonds remains elusive. Negatively charged GeV (S=1/2 ground state) has been investigated using ODMR for quantum memory applications at 300 mK and found to exhibit a 440 s and 1.46 s [43]. These defects were found to be particularly responsive to dynamical decoupling protocols with reaching 24 ms, representing a factor of two improvement compared to negatively charged SiV. SnV (S=1/2 ground state) are also robust spin centres and have been the subject of spin control experiments. Rosenthal et al., report 20 ms and 170 s when measured at 1.7 K in highly strained SnV-centres, which comprises spin lifetime for significant improvement in operation fidelity[44]. Trusheim et al., report a longer s at 1.26 ms and 540 ns at 2.9 K in a less strained system[45]. Negatively charged NiV-centres (S=1/2) are near-infrared emitters, making them especially interesting for quantum communications due to their compatibility with conventional optical cables[26]. The S=1/2 ground state of the negatively charged NiV-centre has a predicted 0.1 ms coherence time at 4 K[46], but only recently have steady-state ODMR studies been reported[38].
@@
Beyond diamond, silicon carbide (SiC) spin systems also show promising optical and coherence properties for quantum applications[47, 48, 49, 50, 51, 52]. SiC is a complex material with over 200 polymorphs. It is also used commercially in electronics and hence benefits from decades of manufacturing experience. Mercifully, the study of quantum systems has largely been restricted to 3C-, 4H-, and 6H-SiC, where C and H signify a cubic and hexagonal structure, respectively, and the preceding number designates its polytype[48]. Pure SiC has a wide bandgap (2 - 3 eV), weak spin-orbit coupling, and a naturally low abundance of nuclear spins. Importantly, it is capable of harbouring several different optically addressable colour centres with (often) near-infrared emission and record ODMR contrasts[53]. The most commonly studied defects consist of neutral or negatively-charged monovacancies (VC or V) and divacancies (VV) but can also include carbon anti-site vacancies (CVC), charged NVs, Cr4+-,[54, 55] V3+, V4+[56], Mo5+[57], and Ti-centres[58, 48, 52]. Further layers of complexity are added by consideration of additional oxidation state and crystallographic sites within each polytype. Understandably, studies often focus on the "best performing" defect within a given sample, and seldom are single samples of SiC homogeneous. This represents the materials most significant practical limitation as a spin qubit candidate compared to other FOAM platforms.
@@
Nevertheless, several defects exhibit outstanding spin-optical properties and the fabrication challenges are beginning to be addressed. Negatively charged V in isotopically purified (28Si) 4H-SiC exhibit the largest ODMR contrast at 97 at 4 K[53]. With a S=3/2 ground state, this defect exhibits ZFS associated of a few MHz with degenerate pairs of <1/2| and <3/2| states. Under a magnetic field (82 mT) precisely aligned to the crystallographic c-axis, this degeneracy is lost and in the excited, so-called "V1" state, <3/2| shift higher energy than the <1/2| states. To achieve 97 contrast, the authors first equilibrate the spin populations using a 40 s off-resonance pump (at 730 nm), followed by an on-resonance pump (at 861 nm) lasting up to 80 s. On-resonance optical pumping causes <3/2| states to selectively decay by into a non-radiative metastable state followed by spin-selective repopulation of <1/2| ground states thereby generating a spin polarisation of up to 90. The authors record and of 0.80.12 ms and 302 s, respectively. More recently, it was shown that V defects can be implanted into nanophotonic waveguides fabricated from 4H-SiC while also controlling the alignment of individual defects and maintaining excellent spin-optical properties[59]. Here, of bulk V-centres was measured at 344 s at 10 K, whilst those in the 1 s diameter waveguides measured at 9.40.7 s. The coherence properties can be further improved by a factor of 10 through a regime of isotopic purification, and another factor 5 by reducing strain inhomogeneity through a regime of annealing. Using this combined approach, Lekavicius et al., enhanced from 400 ns to 20 s at RT[60].
@@
By comparison, divacancies also exhibit compelling spin-optical properties albeit with lower optical contrasts. Divacancies are formed by annealling pre-irradiated SiC at over 700oC[61, 62]. Conversion efficiency into divacancies only reaches a few percent, which may in part be due to counterproductive divacancy-dissociation back into VC, V, and CVC-centres during the annealing process[63]. As S=1 ground state species, divacancies exhibit more intense luminescence from their mS= states compared to their mS=0 state and can optically initialised the same fashion as NV- diamond thanks to its C-symmetry. The kk-divacancy (a neutral VV-centre) in 4H-SiC with natural isotopic abundance can demonstrate times of 1.3 ms at 20 K when decoupled from 13C and 29Si nuclear spins under a 30 mT field[64], significantly longer than NV- diamond under similar conditions. There are also several unknown undefined divacancy-type centres that demonstrate remarkable insensitivity to temperature. For example, Yan et al., reported that PL8-centres, which are triplet ground state species, exhibit similar quantum spin properties at 20 K, with measured and of 15.60.5 s and 18410 ns, and at RT, with measured and of 9.10.1 s and 1809 s[65]. This long-lived RT spin coherence is shared by PL6-centres which are distinguishable by their 1038 nm (vs PL8’s 1007 nm) ZPL emission, demonstrating that several species in a SiC matrix are suitable for RT quantum applications. Moreover, divacancy systems were the first spin-centres demonstrated to be amendable to all electrical spin-ensemble readout and initialisation[66, 67]. These functionalities have since been discovered with monovacancies under ambient conditions[68], demonstrating a strong potential to avoid some of the difficulties associated with ODMR spectroscopy such as photon collection efficiency.

### v2

# Surveying optically addressable spin qubits for quantum information and sensing technology

@@ Abstract
Quantum technologies offer ways to solve certain tasks more quickly, efficiently, and with greater precision than their classical counterparts. Yet substantial challenges remain in the construction of sufficiently error-free and scalable quantum platforms needed to unlock any real benefits to society. Acknowledging that this hardware can take vastly different forms, our review here focuses on so-called spintronic (i.e. spin-electronic) materials that use electronic or nuclear spins to embody qubits. Towards helping the reader to spot trends and pick winners, we have surveyed the various families of optically addressable spin qubits and attempted to benchmark and identify the most promising ones in each. We go on to reveal further trends that demonstrate how qubit lifetimes depend on the material’s synthesis, the concentration/distribution of its embedded qubits, and the experimental conditions.
Keywords spin qubits quantum sensing spintronics materials engineering
@@ 1 Introduction
Over the last 40 years, quantum computation has progressed from concept to hardware. In 2023, IBM reported the commissioning of a 1121-qubit superconducting quantum processor [1], while Honeywell demonstrated a quantum charge-coupled device architecture based on trapped ions [2]. Despite such achievements, the scalability and, thus ultimate utility of each species of hardware remains contested.
@@
Recently, spintronic (spin-electronic) materials have garnered attention as contenders for use as qubit media. The best of these materials offer coherence times exceeding milliseconds, albeit at cryogenic temperatures [3]. Progress in the field has made the dream of usefully storing quantum information and/or implementing quantum processing at scale with appropriately interacting spins more tangible. For example, the start-up company Quantum Brilliance recently demonstrated a room-temperature qubit system based on nitrogen-vacancy (NV) centres in diamond [4]. Beyond computation, these same spintronic materials are already providing advantageous forms of quantum sensing, especially under ambient conditions where thermally-driven electronic and photonic noise present significant challenges for quantum applications [5]. Their advantages stem from the ability to optically initialise spin states into highly non-Boltzmann populations and, likewise, to implement optical read-out routines that are less impacted by blackbody radiation.
@@
In quantum spintronic materials, information is most simply encoded in the spin state of a two-level system, but can be encoded across multiple levels in high-spin (S1) systems, so facilitating error correction [6]. For qubits, a magnetic or electric field is applied to split otherwise degenerate spin states. The higher and lower energy spin states, and , form two orthogonal basis states. A pure state is formed as a coherent superposition of the two (, see Figure 1a). Geometrically, we can visualise any such state as a point on the surface of a Bloch sphere, the location being defined by a polar angle and an azimuthal angle [7]. The value of determines the probability amplitude of finding the qubit in either the or state, while determines the relative phase of the qubit between the and states.
To be useful, spintronic materials must enable both single-spin addressability and long storage times at high levels of fidelity. These properties are, in turn, limited by the materials’ spin-lattice relaxation () and spin decoherence times (, also known as phase memory time). and represent the time it takes for the "longitudinal" and "transverse" magnitude of the spin state vector, respectively, to decay by a factor of e. In practical terms, is the lifetime of a non-Boltzmann population (Figure 1b), and, put simply by DiVincenzo [8], characterises the interactions of a qubit with its environment. These interactions are ultimately responsible for causing a qubit in a generic (pure) state to relax to a mixed state, leading to the loss of quantum information via phase decoherence (Figure 1c). These parameters can be measured using electron spin resonance (ESR) techniques, which measure light/microwave absorptions and emissions from a material as a function of frequency or magnetic field strength (Figure 2). is typically measured by Hahn-Echo pulses but can also be determined from the homogeneous emission linewidth of its resonance frequency, (Figure 1d). In practice, spin systems exhibit an inhomogeneously broadened linewidth, , due to differences in the magnetic field, B, at the locations of otherwise equivalent spins (see Figure 1e). These differences are in turn due to differences in the locations and orientations of the electronic and/or nuclear spins that surround each measured spin. Following Chavhan et al., the relationship between and can be represented as 1/ = 1/ + B, where is the gyromagnetic ratio [9]. In general, sets an upper limit on and, since energy relaxation proceeds more slowly than decoherence, is (often substantially) longer than [10]. Likewise, sets an upper limit on (the pulse sequence used to measure deliberately attempts to reduce the effects of inhomogeneity). Thus, .
Quantum systems capable of operating at room temperature open up many additional applications that the overhead of cryogenic operation precludes. But, achieving high fidelity (in the initialisation, gate operations and read-out) of room-temperature qubits remains extremely challenging. In this review, we have attempted to survey the available "fully optically addressable materials" (FOAMs) that are most directly relevant to quantum information and sensing technology [11]. We do not provide an exhaustive account of all reported spin systems. Rather, we present the different approaches, materials-wise, that have shown promise and report the properties (where available) of the best-performing representatives of each approach.
@@ 2 Overview of Candidate Spin Qubit Materials
FOAMs are an attractive option for quantum applications since they typically involve spin levels that are energetically separated by far more than and can be initialised using visible/near-infrared light to generate quasi-"pure" quantum states. Furthermore, optical signals, being themselves far larger than , are less affected by thermal noise and benefit from excellent single-photon detectors that are available even at room temperature. Since the energies of the photons involved in both processes greatly exceed , room-temperature operation at high levels of fidelity is possible. Such materials are usually probed using ODMR spectroscopy, where the quantum sensing sensitivity for a.c. magnetic fields (for example), , is proportional to , where is the duration of the readout process, is the measurement contrast, is the number of spins and is the number of scan averages. Relatively few FOAMs have been demonstrated so far and are often limited by either low photoluminescence yields or relatively short-lived spin polarizations () and coherences (), which have been summarised in Figure 3 to support our discussion. As a prerequisite for inclusion in figure 3, we have opted to follow the criteria put forward by Weber et al., for the realization of an optically addressable solid-state spin qubit (that is useable)[12]:
@@
2.1 3D spin-defect FOAMs
2.1.1 Diamond spin centres
@@
A well-established example of a FOAM which satisfies the above criteria is the negatively charged nitrogen-vacancy centres in diamond (herein simply, NV-diamond), which utilise the triplet ground state and hyperfine splitting arising from the NV-centre. Initialisation (i.e., electron spin polarisation) is achieved by optical excitation with green light. The newly generated excited states either relax by emission of 637 nm light or undergo intersystem crossing (ISC) into a metastable singlet state (Figure 4a). Repopulation of the ground state then follows an intermediate relaxation between two singlet states (with emission at 1042 nm) and finally spin-selective ISC into the T0 sublevel, resulting in a strong spin polarisation. At zero-applied magnetic field (ZF) the T±1 states are degenerate due to the defects C3V symmetry. The application of a magnetic field lifts this degeneracy, which enables spin manipulation using microwave pulses, thereby modifying the fluorescence at characteristic Zeeman splitting frequencies (), which can be detected following subsequent optical excitations and fluorescence back to T0.
NV-diamond has been a cornerstone of ODMR-based quantum sensing due to its robust spin properties even at room temperature. Due to the relatively low spin densities (resulting in small dipolar coupling between spins) and the mismatch between diamond lattice vibrations (phonons) and the Larmor frequency of electron spins in NV-’s S=1 ground state, NV-diamond can exhibit s of several milliseconds at room temperature [60, 61]. The nature and mechanism of its spin-lattice relaxation as a function of temperature, which includes an ‘Orbach-like" process (dependent on the phonon density at the spin transition frequency) and spin-phonon Raman scattering (of either first [62] or else second [61] order) scaling like , have been repeatedly investigated.
@@
The material’s quantum spin properties are highly dependent on the NV-centre depth [63], concentration [64], crystal strain [65], and the presence of impurities such as 13C, N-centres or EPR-inactive neutral or positively charged NV-centres [66, 67]. Its popularity has seen a plethora of investigations to understand and modulate its spin properties. For example, high-field EPR spectroscopy and temperature-dependent measurements have demonstrated that decoherence from the N and 13C flip-flop fluctuations can be almost eliminated at low temperatures where the spin-bath is polarised [68]. Here, reached 250 s at 2 K, following a sharp increase below 12 K in high temperature high pressure (HTHP) diamond samples. Achieving similar properties under low-field and higher temperature conditions with thermally polarised nuclei is challenging and requires meticulous materials preparation and/or the use of dynamical decoupling methods. For example, the impact of parasitic nuclear spins and impurities was shown most remarkably by Balasuramanian et al., [22]. Careful growth by CVD on a diamond substrate using isotopically enriched feedstock led to just 0.3 13C abundance and low levels of other paramagnetic impurity, resulting in 1.8 ms at room-temperature.
@@
More recently, increasing attention has been paid to using diamonds as a host for other spin defects owing to diamond’s wide band gap, efficient thermal dissipation, and physical and chemical stability. As such, various magnetically-active colour centres have been investigated for quantum applications [69]. These include HV [70], BV [71], OV [72], so-called group-IV vacancy defects [73] such as SiV [74, 75], SnV [76, 57], GeV [77], and PbV [78, 79], and even transition metal defects originating as impurities like NiV [80]. To our knowledge, spin polarisation has not been observed in HV or OV centres, whilst ODMR experiments on BV and PbV have not been reported.
Metal-vacancy spin-centres typically exhibit strong spin-orbit coupling (several hundred GHz) that gives rise to zero-field splitting (ZFS) even for S=1/2 species, as well as electron-nuclear spin coupling, making them qudit candidates (Figure 4b-d). Moreover, group-IV defects otherwise exhibit useful photonic properties that make them attractive for quantum applications such as Fourier-limited zero-phonon line (ZPL) linewidths, high spectral stability, coherent photon emission, and strain-responsive band gap engineering [81].
For example, negatively charged SiV (S=1/2 system, (Figure 4c) has enabled direct observation of photon interference [82]. Outside of a dilution refrigerator, the spin coherence lifetimes are severely impacted by thermal acoustic phonon coupling. At 100 mK, reaches 1 second while the longest measured was 1.5 s. Spin coherence could be maintained by dynamical decoupling up to 600 mK where the coherence time during a dynamical decoupling scheme () measured 60 s. can also be improved through strain engineering, which modifies the spin-orbit coupling, which is responsible for inducing the ground state splitting, and subsequently the spin-phonon coupling [83]. By comparison, the neutral SiV (S=1 ground state) demonstrates an impressive 25 s, 0.1 ms even at 15 K, decreasing to 7.8 and 2 s at room temperature [23]. However, a route to reliably synthesising SiVs in diamonds remains elusive.
Negatively charged GeV (S=1/2 ground state, (Figure 4d) has been investigated using ODMR for quantum memory applications at 300 mK and found to exhibit a 440 s and 1.46 s [26]. These defects were found to be particularly responsive to dynamical decoupling protocols with reaching 24 ms, representing a factor of two improvement compared to negatively charged SiV.
SnV-s (S=1/2 ground state) are also robust spin centres and have been the subject of spin control experiments. Rosenthal et al., report 20 ms and 170 s when measured at 1.7 K in highly strained SnV-centres, which comprises spin lifetime for significant improvement in operation fidelity [25]. Trusheim et al., report a longer 1.26 ms and 540 ns at 2.9 K in a less strained system [58].
Negatively charged NiV-centres (S=1/2) are near-infrared emitters, making them especially interesting for quantum communications due to their compatibility with conventional optical cables [69]. The S=1/2 ground state of the negatively charged NiV-centre has a predicted 0.1 ms coherence time at 4 K [84], but only recently have steady-state ODMR studies been reported [80].
2.1.2 Silicon Carbide Spin Centres
@@
Beyond diamond, silicon carbide (SiC) also shows promising optical and coherence properties for quantum applications [85, 86, 87, 88, 89, 90]. SiC is a complex material with over 200 polymorphs. It is also used commercially in electronics and hence benefits from decades of manufacturing experience. Mercifully, the study of quantum systems has largely been restricted to 3C-, 4H-, and 6H-SiC, where C and H signify the cubic and hexagonal structures, respectively, and the preceding number designates its polytype [86] (see Figure 5a). Pure SiC has a wide bandgap (2 - 3 eV), weak spin-orbit coupling, and a naturally low abundance of nuclear spins. Importantly, it is capable of harbouring several different optically addressable colour centres with (often) near-infrared emission and record ODMR contrasts[31]. The most commonly studied defects consist of neutral or negatively-charged monovacancies (VC or VSi) and divacancies (VV) but can also include carbon anti-site vacancies (CSiVC) (see Figure 5b and c), charged NVs, Cr4+- [41, 91], V3+, V4+ [92], Mo5+ [44], and Ti-centres [93, 86, 90]. Further layers of complexity are added by consideration of additional oxidation states and crystallographic sites within each polytype. Understandably, studies often focus on the "best performing" defect within a given sample, and seldom are single samples of SiC homogeneous. This represents the material’s most significant practical limitation as a spin qubit candidate compared to other platforms.
@@
Nevertheless, several defects exhibit outstanding spin-optical properties, and the fabrication challenges are beginning to be addressed. Negatively charged VSi in isotopically purified (28Si) 4H-SiC exhibits the largest ODMR contrast at 97 at 4 K [31]. With an S=3/2 ground state, its associated ZFS is a few MHz with degenerate pairs of |1/2 and |3/2 states (Figure 5d). Under a magnetic field (82 mT) precisely aligned to the crystallographic c-axis, this degeneracy is lost and in the excited, so-called "V1" state, |3/2 shifts higher energy than the |1/2 states. To achieve 97 contrast, the authors first equilibrate the spin populations using a 40 s off-resonance pump (at 730 nm), followed by an on-resonance pump (at 861 nm) lasting up to 80 s. On-resonance optical pumping causes |3/2 states to selectively decay into a non-radiative metastable state, followed by spin-selective repopulation of |1/2 ground states. This leads to a spin polarisation of up to 90 which is paired with record s and s of 0.80.12 ms and 302 s, respectively. More recently, it was shown that VSi defects can be implanted into nanophotonic waveguides fabricated from 4H-SiC while also controlling the alignment of individual defects and maintaining excellent spin-optical properties [32]. Here, of bulk VSi-centres was measured at 344 s at 10 K, whilst those in the 1 m diameter waveguides measured at 9.40.7 s. The coherence properties can be further improved by a factor of 10 through a regime of isotopic purification, and another factor of 5 by reducing strain inhomogeneity through a regime of annealing. Using this combined approach, Lekavicius et al., enhanced from 400 ns to 20 s at room-temperature [33].
@@
By comparison, divacancies also exhibit compelling spin-optical properties, albeit with lower optical contrasts. Divacancies are formed by annealing pre-irradiated SiC at over 700oC [97, 98]. Conversion efficiency into divacancies only reaches a few percent, which may in part be due to counterproductive divacancy-dissociation back into VC, VSi, and CSiVC-centres during the annealing process [99]. As S=1 ground state species, divacancies exhibit more intense luminescence from their mS= states compared to their mS=0 state and can be optically initialised in the same fashion as NV-diamond thanks to its C3V-symmetry. The kk-divacancy (a neutral VSiV-centre) in 4H-SiC with natural isotopic abundance can demonstrate times of 1.3 ms at 20 K when decoupled from 13C and 29Si nuclear spins under a 30 mT field [37], significantly longer than NV-diamond under similar conditions. There are also several divacancy-type centres that demonstrate remarkable insensitivity to temperature. For example, Yan et al., reported that PL8-centres, which are triplet ground state species, exhibit similar quantum spin properties at 20 K, with measured and of 15.60.5 s and 18410 ns, and at room-temperature, with measured and of 9.10.1 s and 1809 s [36] (Figure 5e-g). This long-lived room-temperature spin coherence is shared to a reduced extent by PL6-centres, which are distinguishable by their 1038 nm (vs PL8’s 1007 nm) ZPL emission, demonstrating that several species in a SiC matrix are suitable for room-temperature quantum applications. Moreover, divacancy systems were the first spin-centres demonstrated to be amenable to an all-electrical spin-ensemble readout and initialisation scheme [100, 101]. These functionalities have since been discovered with monovacancies under ambient conditions [102], demonstrating a strong potential to avoid some of the difficulties associated with ODMR spectroscopy, such as pump-light stability and photon collection efficiency.
2.1.3 Inorganic Semiconductor Spin Centres